OpenStudy (anonymous):

Find the solution of this system of equations algebraically. (A) y = -x (B) y = x3 + 3x2 + 2x (0, 0) (0, -1) and (0, -2) (0, 0), (-1, 1) and (-2, 2) (-1, 1) and (-3, 3) It has no solution.

OpenStudy (anonymous):

I get 0,-1,-2 for the x values

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

hey

OpenStudy (anonymous):

hey

OpenStudy (freckles):

did you replace y in the equation with -x that is did you solve: \[-x=x^3+3x^2+2x\]

OpenStudy (anonymous):

I got x to equal 0 on the equation you just put

OpenStudy (freckles):

you add x on both sides \[0=x^3+3x^2+3x \\0=x(x^2+3x+3)\] and yes one solution is x=0 but what happens when you solve x^2+3x+3=0?

OpenStudy (anonymous):

I get 0 still

OpenStudy (freckles):

well yes the only real solution you get from the equation 0=x(x^2+3x+3) is zero because x^2+3x+3=0 gives you complex solutions

OpenStudy (anonymous):

so the answer is a?

OpenStudy (freckles):

so (0,0) is the only point of intersection